the sum of the first and the third terms of a geometric progression is 40. the sum of the second and the fourth terms is 120. Determine the first term and the common ration of the progression.
Answer by mananth(16946)

let the terms be a,ar,ar^2,ar^3
First condition
a+ar^2=40
a(1+r^2)=40
(1+r^2)=40/a
II condition
ar+ar^3=120
ar(1+r^2)=120
substitute (1+r^2)
ar(*(40/a) = 120
40r= 120
r=3
a(1+r^2)= 40
substitute r
a(1+9)=40
10a=40
a=4
CHECK
4, 12, 36,108,